A mechanic applies a torque to a nut with a wrench 0.250 m long. Because of the cramped space, he must exert a force upward at an angle of 50.0° with respect to a line from the nut through the end of the wrench. If the force he exerts has magnitude 40.0 N, what magnitude torque (in N · m) does he apply to the nut? (Assume that the force is applied at the end of the wrench.)



___ N · m



1
Expert's answer
2021-10-11T08:55:37-0400

We can find the torque as follows:


"\tau = Frsin\theta=40.0\ N\cdot0.250\ m\cdot sin50.0^{\circ}=7.66\ N\cdot m."